Zeronknight (zeronknight):

Solve the following inequality: −2x−12≥36 A. x\leq -24 B. x\geq -6 C. x\leq -6 D. x\geq -24

Zeronknight (zeronknight):

\[A. x\leq -24\]

Zeronknight (zeronknight):

\[B. x\geq -6 \]

Zeronknight (zeronknight):

\[C. x\leq -6\]

Zeronknight (zeronknight):

\[D. x\geq -24 \]

OpenStudy (anonymous):

solving rn. hold on

Zeronknight (zeronknight):

You would add 12 then divide by 2, but why is 24 negative?

OpenStudy (anonymous):

You would divide by a -2 and a negative divided by a positive results in a negative answer

Zeronknight (zeronknight):

ohh

OpenStudy (anonymous):

So it's D

Zeronknight (zeronknight):

ye, thank you for clarifying that! Welcome to OpenStudy! :)

OpenStudy (anonymous):

Your Welcome :p and Thank you hahaha